# Learning Prompt with Distribution-Based Feature Replay for Few-Shot Class-Incremental Learning

### v1

@@ Abstract
Few-shot Class-Incremental Learning (FSCIL) aims to continuously learn new classes based on very limited training data without forgetting the old ones encountered. Existing studies solely relied on pure visual networks, while in this paper we solved FSCIL by leveraging the Vision-Language model (e.g., CLIP) and propose a simple yet effective framework, named Learning Prompt with Distribution-based Feature Replay (LP-DiF). We observe that simply using CLIP for zero-shot evaluation can substantially outperform the most influential methods. Then, prompt tuning technique is involved to further improve its adaptation ability, allowing the model to continually capture specific knowledge from each session. To prevent the learnable prompt from forgetting old knowledge in the new session, we propose a pseudo-feature replay approach. Specifically, we preserve the old knowledge of each class by maintaining a feature-level Gaussian distribution with a diagonal covariance matrix, which is estimated by the image features of training images and synthesized features generated from a VAE. When progressing to a new session, pseudo-features are sampled from old-class distributions combined with training images of the current session to optimize the prompt, thus enabling the model to learn new knowledge while retaining old knowledge. Experiments on three prevalent benchmarks, i.e., CIFAR100, mini-ImageNet, CUB-200, and two more challenging benchmarks, i.e. SUN-397 and $*{}^{*}$ proposed in this paper showcase the superiority of LP-DiF, achieving new state-of-the-art (SOTA) in FSCIL. Code is publicly available at https://github.com/1170300714/LP-DiF.
1 Introduction
@@
    Extensive evaluations and comparisons on three prevalent FSCIL benchmarks (CIFAR-100, CUB-200 and mini-ImageNet) and two proposed more challenging benchmarks (SUN-397 and $*{}^{*}$) show the superiority of our methods in comparison to state-of-the-arts.
2 Related Work
@@
Few-Shot Class-Incremental Learning. The few-shot class-incremental learning methods (FSCIL) aims to train a model in a class-incremental manner [10, 59] with only a few samples for each new tasks [39]. Existing studies can be categorized into four families, i.e., dynamic network-based methods, meta-learning-based methods, feature space-based methods, and replay-based methods. In specific, dynamic network structure [37, 49, 50, 15] is proposed to adaptive learn the new knowledge. Meta learning-based methods [28, 61, 18, 9, 53, 63] employ a session sampling scheme, simulating the incremental learning process during evaluation. feature space-based methods [8, 2, 23, 55, 57, 58, 62], focus on mapping the original image into a condensed feature space while preserving its essential attributes. Replay-based methods [26, 7, 11] retain or produce significant data from prior tasks to be reintroduced in the ongoing task. While these methods have shown commendable performance, all those studies are based on feature extractors and classifiers built from deep networks trained in the base session. Due to the scarcity of incremental class samples, the feature representation ability is limited. In contrast, we propose to construct an incremental learner on a VL pre-trained model [29, 60] that offers inherent merits for FSCIL, i.e., endowing the image encoder with powerful feature representation abilities.
@@
Replay-based Incremental Learning. The replay-based approach in incremental learning leverages knowledge from previous tasks to mitigate catastrophic forgetting in models [5, 4, 3, 20, 31, 6, 34, 16, 19, 1]. A basic data replay approach involves retaining a concise exemplar set, capturing essential samples from prior tasks [5, 4], then, the classification model is trained on the combination of exemplars and the data of the current task. Different from directly storing the real instances, several following works [34, 16, 19, 22] leveraged a generative model [14, 24] for generating data from previous tasks. Compared to methods based on real image replay, pseudo replay reduces storage needs by eliminating the requirement for exemplars and enriches the diversity of samples from previous tasks. Yet, the overhead of training the image generator and dynamically producing pseudo images introduces additional computational demands and prolongs training time. Instead of retaining an image generator, we represent the feature representation for each class using a Gaussian distribution, utilizing it to sample pseudo-features for rehearsing prior knowledge. Moreover, drawing samples from this distribution is computationally efficient, offering our method an effective way for handling prevent catastrophic forgetting.
Incremental Learning via Pre-trained Model. Recent studies have explored constructing incremental learners using pre-trained models [13, 38, 46, 45, 35, 43, 44, 52]. The core idea of these methods is to leverage a pre-trained backbone, e.g., ViT [12], for robust image feature extraction, while only fine-tuning a selected set of parameters to adapt to new tasks. For example, L2P [46] employs a fixed pre-trained ViT as its backbone and sustains a dynamic prompt pool with various sets of adaptable prompts. Some following works [43, 35] built upon this concept, applying it to the VL pretrained model [29], leveraging linguistic knowledge to bolster classification performance. The above studies underscore the significant advantages of using pretrained models to boost performance in standard CIL scenarios. As for FSCIL, we inherit the advantages of pretrained models in CIL, and further maintain a Gaussian distribution at the feature level for each class to mitigate catastrophic forgetting.
@@
3 Proposed Method
@@ 3.1 Approach Overview
However, using only the $\mathcal{D}^{(t)}_{\text{Train}}$ to optimize the prompt in session $t$ will inevitably lead to catastrophic forgetting. Ideally, learning prompt with all training data from both previous and current sessions (e.g. $\bigcup^{t}_{s=0}\mathcal{D}^{(s)}_{\text{Train}}$) can address this issue, but this is not allowed under the protocol of FSCIL. Therefore, this paper adopts a compromise solution, proposing to record old knowledge by maintaining statistical distributions of old classes instead of directly storing origin images. We setup a feature-level Gaussian distribution to represent each old class, which is represented by a mean vector and a diagonal covariance matrix. We name it the old-class distribution. The mean vector and diagonal covariance matrix of the old-class distribution are estimated jointly from the features of real images as well as synthetic features generated by a VAE decoder. When learning the prompt in a new session, we randomly sample features based on the statistical distribution of old classes to replay old knowledge. Then, the sampled features of old classes and the real features of new classes will jointly optimize the prompt, thereby learning new knowledge while also replaying old knowledge. In the following, We will introduce how to obtain the old-class distribution in Sec 3.2, and how to learn prompt in Sec 3.3.
3.2 Estimation of Old-Class Distribution
@@
$\mathcal{V}_{\text{VAE}}(\mathbf{z})=\{[\mathbf{V}_{\text{VAE}}]_{l}+\bm{r}\}^{L}_{l=1},$ (5)
@@
3.3 Learning Prompt with Feature Replay
@@
    When $t=0$, i.e., the first session, we just follow the approach in Sec.3.1, randomly initializing $\mathcal{V}$ and learning them with $\mathcal{D}^{(0)}_{\text{Train}}$ by $\mathcal{L}_{\text{n}}$ (Eq. (3)).
@@
4 Experiments
4.1 Experimental Setup
Table 1: Comparison on mini-ImageNet, where the rows filled with gray indicate the existing SOTA FSCIL methods. “Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance.
Methods Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
@@ 4.1 Experimental Setup
TOPIC [37] $61.31$ $50.09$ $45.17$ $41.16$ $37.48$ $35.52$ $32.19$ $29.46$ $24.42$ $39.64$ $36.89$
CEC [54] $72.00$ $66.83$ $62.97$ $59.43$ $56.70$ $53.73$ $51.19$ $49.24$ $47.63$ $57.75$ $24.37$
F2M [33] $72.05$ $67.47$ $63.16$ $59.70$ $56.71$ $53.77$ $51.11$ $49.21$ $47.84$ $57.89$ $24.21$
Replay [27] $71.84$ $67.12$ $63.21$ $59.77$ $57.01$ $53.95$ $51.55$ $49.52$ $48.21$ $58.02$ $23.63$
MetaFSCIL [9] $72.04$ $67.94$ $63.77$ $60.29$ $57.58$ $55.16$ $52.90$ $50.79$ $49.19$ $58.85$ $22.85$
GKEAL [62] $73.59$ $68.90$ $65.33$ $62.29$ $59.39$ $56.70$ $54.20$ $52.59$ $51.31$ $60.48$ $22.28$
FACT [57] $72.56$ $69.63$ $66.38$ $62.77$ $60.60$ $57.33$ $54.34$ $52.16$ $50.49$ $60.70$ $22.07$
C-FSCIL [18] $76.40$ $71.14$ $66.46$ $63.29$ $60.42$ $57.46$ $54.78$ $53.11$ $51.41$ $61.59$ $14.99$
BiDistFSCIL [56] $74.65$ $70.70$ $66.81$ $63.63$ $61.36$ $58.14$ $55.59$ $54.23$ $53.39$ $62.06$ $21.26$
FCIL [15] $76.34$ $71.40$ $67.10$ $64.08$ $61.30$ $58.51$ $55.72$ $54.08$ $52.76$ $62.37$ $23.58$
SAVC [36] $81.12$ $76.14$ $72.43$ $68.92$ $66.48$ $62.95$ $59.92$ $58.39$ $57.11$ $67.05$ $24.01$
NC-FSCIL [51] $84.02$ $76.80$ $72.00$ $67.83$ $66.35$ $64.04$ $61.46$ $59.54$ $58.31$ $67.82$ $25.71$
CLIP (Baseline[29] $80.01$ $79.16$ $78.89$ $77.97$ $77.44$ $76.83$ $76.32$ $76.02$ $75.45$ $77.57$ 4.56
LP-DiF (Ours) 96.34 96.14 94.62 94.37 94.06 93.44 92.21 92.29 91.68 93.76 $4.66$
Datasets. Following the mainstream benchmark settings [56], we conduct experiments on three datasets, i.e., CIFAR-100 [25], mini-ImageNet [32] and CUB-200 [41], to evaluate our LP-DiF. Additionally, this paper also proposes two more challenging benchmarks for FSCIL, i.e., SUN-397 [48] and $*{}^{*}$. For SUN-397 [48], it has about twice the number of classes and incremental sessions than the aforementioned three benchmarks. We evaluate our method on this benchmark to reveal whether it works in scenarios with more classes and more incremental sessions. $*{}^{*}$ is a variant of CUB-200 but excluding the base session, is used to evaluate whether our method works in scenarios without the base session. Refer to Suppl. for the details of all selected benchmarks.
@@
Implementation Details. All experiments are conducted with PyTorch on 8×8$\times$8 × NVIDIA RTX $2080$Ti GPUs. We leverage the ViT-B/16 as the image encoder and adopt SGD with $0.9$ momentum to optimize the prompts. The learning rate is initialized by $0.002$. For the base session, the batch size is set to $64$ and the training epoch is set to $200$, As for each incremental session, the batch size and the training epochs are set to $25$, $100$, respectively. The VAE component is enabled only for incremental sessions. For the hyper-parameters, $M$ is set to $10$; $B$ and $\lambda_{\text{o}}$ are set to $8$ and $2$, respectively; $L$ is set to $16$ following Zhou et al. [60]; $\lambda_{\text{r}}$ is set to $1$ following Wang et al. [47].
Table 2: Comparison with Upper bound on the three common benchmarks in terms of Avg.

### v3

@@ Abstract
Few-shot Class-Incremental Learning (FSCIL) aims to learn new classes with few examples while retaining knowledge of previously encountered ones. Existing studies relied on pure visual networks, while in this paper we solved FSCIL by leveraging the pretrained vision-language model and propose a simple yet effective framework, named Learning Prompt with Distribution-based Feature Replay (LP-DiF). We observe that using CLIP for zero-shot evaluation significantly outperforms leading methods. Then, prompt tuning is involved to further improve its adaptation ability, enabling continuous learning of session-specific knowledge. To prevent the learnable prompt from forgetting old knowledge, we propose a pseudo-feature replay approach. Specifically, we preserve old knowledge of each class by maintaining a feature-level Gaussian distribution with a diagonal covariance matrix, which is estimated by the features of training images and synthesized features generated from a VAE. When progressing to a new session, pseudo-features are sampled from old-class distributions combined with training images of the current session to optimize the prompt, thus enabling the model to learn new knowledge while retaining old knowledge. Experiments on prevalent benchmarks, i.e., CIFAR100, mini-ImageNet, CUB-200, and more challenging benchmarks, i.e. SUN-397 and CUB-200 proposed in this paper showcase the superiority of LP-DiF, achieving new state-of-the-art (SOTA) in FSCIL. Code is publicly available at https://github.com/1170300714/LP-DiF.
Index Terms:
Few-shot class-incremental learning, continual learning, prompt tunning.
I Introduction
@@
    Extensive evaluations and comparisons on three prevalent FSCIL benchmarks (CIFAR-100, CUB-200 and mini-ImageNet) and two proposed more challenging benchmarks (SUN-397 and CUB-200) show the superiority of our methods in comparison to state-of-the-arts.
II Related Work
@@
Few-Shot Class-Incremental Learning. The few-shot class-incremental learning methods (FSCIL) aims to train a model in a class-incremental manner [13, 62] with only a few samples for each new tasks [42]. Existing studies can be categorized into four families, i.e., dynamic network-based methods, meta-learning-based methods, feature space-based methods, and replay-based methods. In specific, dynamic network structure [40, 51, 52, 18] is proposed to adaptive learn the new knowledge by dynamically expanding the network structure, so that the new knowledge is preserved by the new network structure. Meta learning-based methods [31, 65, 21, 12, 55, 67, 58] employ a session sampling scheme, where a sequence of sessions are sampled from the base session, aiming to mimic the incremental learning process during evaluation, to allows the model to learn how to retain old knowledge under the condition of a small number of new data samples. Feature space-based methods [11, 4, 26, 59, 61, 63, 66, 2, 3], focus on mapping the original image into a condensed feature space while preserving its essential attributes, which ensures that the representations of old category data are not disrupted when the model is trained on new data. Replay-based methods [29, 10, 14] retain or produce significant data from prior tasks to be reintroduced in the ongoing task. These methods are dedicated to selecting the most representative samples of old categories, or utilizing generative models to produce high-quality pseudo-samples of old categories. While these methods have shown commendable performance, all those studies are based on feature extractors and classifiers built from deep networks trained in the base session. Due to the scarcity of incremental class samples, the feature representation ability is limited. In contrast, we propose to construct an incremental learner on a VL pre-trained model [32, 64] that offers inherent merits for FSCIL, i.e., endowing the image encoder with powerful feature representation abilities.
@@
Replay-based Incremental Learning. The replay-based approach in incremental learning leverages knowledge from previous tasks to mitigate catastrophic forgetting in models [7, 6, 5, 23, 34, 8, 37, 19, 22, 1, 9]. A basic data replay approach involves retaining a concise exemplar set, capturing essential samples from prior tasks [7, 6], then, the classification model is trained on the combination of exemplars and the data of the current task. Different from directly storing the real instances, several following works [37, 19, 22, 25] leveraged a generative model [17, 27] for generating data from previous tasks. Compared to methods based on real image replay, pseudo replay reduces storage needs by eliminating the requirement for exemplars and enriches the diversity of samples from previous tasks. Yet, the overhead of training the image generator and dynamically producing pseudo images introduces additional computational demands and prolongs training time. Instead of retaining an image generator, we represent the feature representation for each class using a Gaussian distribution, utilizing it to sample pseudo-features for rehearsing prior knowledge. Moreover, drawing samples from this distribution is computationally efficient, offering our method an effective way for handling prevent catastrophic forgetting.
@@
Incremental Learning via Pre-trained Model. Recent studies have explored constructing incremental learners using pre-trained models [16, 41, 49, 48, 38, 45, 46, 53, 57]. The core idea of these methods is to leverage a pre-trained backbone, e.g., ViT [15], for robust image feature extraction, while only fine-tuning a selected set of parameters to adapt to new tasks. For example, L2P [49] employs a fixed pre-trained ViT as its backbone and sustains a dynamic prompt pool with various sets of adaptable prompts. Some following works [45, 38] built upon this concept, applying it to the VL pretrained model [32], leveraging linguistic knowledge to bolster classification performance. In addition, Yang. et al. [53] built a Bayesian model based on a fixed feature extracted by a pretrained backbone. During the test stage, they use this Bayesian model as the classifier to mitigate the forgetting problem. The above studies underscore the significant advantages of using pretrained models to boost performance in standard CIL scenarios. As for FSCIL, we inherit the advantages of pretrained models in CIL. Inspired by Yang. et al. [53], we further explore the potential characteristics of fixed feature, maintain a feature-level Gaussian distribution for each class to preserve the old knowledge, and use it to generate pseudo features to mitigate the catastrophic forgetting.
III Proposed Method
@@ III-A Approach Overview
However, using only the $\mathcal{D}^{(t)}_{\text{Train}}$ to optimize the prompt in session $t$ will inevitably lead to catastrophic forgetting. Ideally, learning prompt with all training data from both previous and current sessions (e.g. $\bigcup^{t}_{s=0}\mathcal{D}^{(s)}_{\text{Train}}$) can address this issue, but this is not allowed under the protocol of FSCIL. Therefore, this paper adopts a compromise solution, proposing to record old knowledge by maintaining statistical distributions of old classes instead of directly storing origin images. We setup a feature-level Gaussian distribution to represent each old class, which is represented by a mean vector and a diagonal covariance matrix. We name it the old-class distribution. The mean vector and diagonal covariance matrix of the old-class distribution are estimated jointly from the features of real images as well as synthetic features generated by a VAE decoder. When learning the prompt in a new session, we randomly sample features based on the statistical distribution of old classes to replay old knowledge. Then, the sampled features of old classes and the real features of new classes will jointly optimize the prompt, thereby learning new knowledge while also replaying old knowledge. In the following, We will introduce how to obtain the old-class distribution in Sec III-B, and how to learn prompt in Sec III-C.
III-B Estimation of Old-Class Distribution
@@
$\mathcal{V}_{\text{VAE}}(\mathbf{z})=\{[\mathbf{V}_{\text{VAE}}]_{l}+\bm{r}\}^{L}_{l=1}.$ (5)
@@
III-C Learning Prompt with Feature Replay
@@
    When $t=0$, i.e., the first session, we just follow the approach in Sec.III-A, randomly initializing $\mathcal{V}$ and learning them with $\mathcal{D}^{(0)}_{\text{Train}}$ by $\mathcal{L}_{\text{n}}$ (Eq. (3)).
@@
IV Experiments
TABLE I: Comparison on mini-ImageNet. “Backbone.” represents the backbone of visual model. “Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance. “bs” is the abbreviation of “baseline”. “UB.” is the abbreviation of “Upper bound”.
Methods Backbone Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
@@ IV Experiments
TOPIC [40] Res18 $61.31$ $50.09$ $45.17$ $41.16$ $37.48$ $35.52$ $32.19$ $29.46$ $24.42$ $39.64$ $36.89$
CEC [56] Res18 $72.00$ $66.83$ $62.97$ $59.43$ $56.70$ $53.73$ $51.19$ $49.24$ $47.63$ $57.75$ $24.37$
F2M [36] Res18 $72.05$ $67.47$ $63.16$ $59.70$ $56.71$ $53.77$ $51.11$ $49.21$ $47.84$ $57.89$ $24.21$
Replay [30] Res18 $71.84$ $67.12$ $63.21$ $59.77$ $57.01$ $53.95$ $51.55$ $49.52$ $48.21$ $58.02$ $23.63$
MF [12] Res18 $72.04$ $67.94$ $63.77$ $60.29$ $57.58$ $55.16$ $52.90$ $50.79$ $49.19$ $58.85$ $22.85$
GKEAL [66] Res18 $73.59$ $68.90$ $65.33$ $62.29$ $59.39$ $56.70$ $54.20$ $52.59$ $51.31$ $60.48$ $22.28$
FACT [61] Res18 $72.56$ $69.63$ $66.38$ $62.77$ $60.60$ $57.33$ $54.34$ $52.16$ $50.49$ $60.70$ $22.07$
C-FSCIL [21] Res12 $76.40$ $71.14$ $66.46$ $63.29$ $60.42$ $57.46$ $54.78$ $53.11$ $51.41$ $61.59$ $14.99$
BDF [60] Res18 $74.65$ $70.70$ $66.81$ $63.63$ $61.36$ $58.14$ $55.59$ $54.23$ $53.39$ $62.06$ $21.26$
FCIL [18] Res18 $76.34$ $71.40$ $67.10$ $64.08$ $61.30$ $58.51$ $55.72$ $54.08$ $52.76$ $62.37$ $23.58$
SAVC [39] Res18 $81.12$ $76.14$ $72.43$ $68.92$ $66.48$ $62.95$ $59.92$ $58.39$ $57.11$ $67.05$ $24.01$
NC-FSCIL [54] Res18 $84.02$ $76.80$ $72.00$ $67.83$ $66.35$ $64.04$ $61.46$ $59.54$ $58.31$ $67.82$ $25.71$
CLIP (Bs.[32] ViT-B/16 $80.01$ $79.16$ $78.89$ $77.97$ $77.44$ $76.83$ $76.32$ $76.02$ $75.45$ $77.57$ 4.56
LP-DiF (Ours) ViT-B/16 96.34 96.14 94.62 94.37 94.06 93.44 92.21 92.29 91.68 93.76 $4.66$
Joint-LP (UB. of ours) ViT-B/16 96.3496.34{{96.34}}96.34 96.0796.07{{96.07}}96.07 95.7595.75{{95.75}}95.75 94.9394.93{{94.93}}94.93 94.6194.61{{94.61}}94.61 94.2694.26{{94.26}}94.26 93.9993.99{{93.99}}93.99 93.8393.83{{93.83}}93.83 93.5693.56{{93.56}}93.56 94.8194.81{{94.81}}94.81 2.782.78{{2.78}}2.78
TABLE II: Details of selected benchmarks. The first three lines are commonly used benchmarks, while the last two lines are the more challenging benchmarks proposed in this paper. $\bm{|\mathcal{C}^{\text{All}}|}$, $\bm{|\mathcal{C}^{\text{Base}}|}$ and $\bm{|\mathcal{C}^{\text{Inc}}|}$ denotes the total number of classes, the number of classes in base session, and the number of classes in each incremental session respectively. #Base and #Inc denote the number of base sessions and the incremental session respectively. Shot denotes the number of training images of each incremental session. represents a variant version.
CIFAR-100 [28] 100 60 5 1 8 5
mini-ImageNet [35] 100 60 5 1 8 5
CUB-200 [43] 200 100 10 1 10 5
SUN-397 [50] 397 197 10 1 20 5
CUB-200* [43] 200 0 10 0 20 5
IV-A Datasets and Metrics
Datasets. Following the mainstream benchmark settings [60], we conduct experiments on three datasets, i.e., CIFAR-100 [28], mini-ImageNet [35] and CUB-200 [43], to evaluate our LP-DiF. Tab. II summarizes the details of each selected benchmark.
Additionally, this paper also proposes two more challenging benchmarks for FSCIL, i.e., SUN-397 [50] and CUB-200.
  • CUB-200 is a variant of CUB-200 but excludes the base session. We evenly divide the total $200$ classes into $20$ incremental sessions, with each session containing $10$ categories. Following the previous study [60], there are $100$ classes in the base session, and the remaining classes will be divided into $10$ incremental sessions, with each incremental session comprising $10$ classes. We use it to evaluate whether our method works in scenarios without the base session.
Metrics. Following existing FSCIL methods [60, 39, 40], we employ the Avg., which is the average accuracy of each session, as primary metric for performance comparison. In addition, we also employ the performance drop rate (PD.), which represents the drop of performance of the last session compared to the first session, to reflect the extent of the model’s forgetting of old knowledge.
IV-B Implementation Details.
@@ IV-A Datasets and Metrics
All experiments are conducted with PyTorch on 8×8$\times$8 × NVIDIA RTX $2080$Ti GPUs. We leverage the ViT-B/16 as the image encoder of LP-DiF and adopt SGD with $0.9$ momentum to optimize the prompts. The learning rate is initialized by $0.002$. For the base session, the batch size is set to $64$ and the training epoch is set to $200$, As for each incremental session, the batch size and the training epochs are set to $25$, $100$, respectively. The VAE component is enabled only for incremental sessions. For the hyper-parameters, $M$ is set to $10$; $B$ and $\lambda_{\text{o}}$ are set to $8$ and $2$, respectively; $L$ is set to $16$ following Zhou et al. [64]; $\lambda_{\text{r}}$ is set to $1$ following Wang et al. [47].
TABLE III: Comparison with state-of-the-art FSCIL methods on CUB-200. “Backbone.” represents the backbone of visual model. “Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance. “bs” is the abbreviation of “baseline”. “UB.” is the abbreviation of “Upper bound”.
Methods Backbone Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
0 1 2 3 4 5  6 7 8 9 10
TOPIC [40] Res18 $68.68$ $62.49$ $54.81$ $49.99$ $45.25$ $41.40$ $38.35$ $35.36$ $32.22$ $28.31$ $26.26$ $43.92$ $42.42$
CEC [56] Res18 $75.85$ $71.94$ $68.50$ $63.50$ $62.43$ $58.27$ $57.73$ $55.81$ $54.83$ $53.52$ $52.28$ $61.33$ $23.57$
Replay [30] Res18 $75.90$ $72.14$ $68.64$ $63.76$ $62.58$ $59.11$ $57.82$ $55.89$ $54.92$ $53.58$ $52.39$ $61.52$ $23.51$
MetaFSCIL [12] Res18 $75.90$ $72.41$ $68.78$ $64.78$ $62.96$ $59.99$ $58.30$ $56.85$ $54.78$ $53.82$ $52.64$ $61.93$ $23.26$
FACT [61] Res18 $75.90$ $73.23$ $70.84$ $66.13$ $65.56$ $62.15$ $61.74$ $59.83$ $58.41$ $57.89$ $56.94$ $64.42$ $18.96$
FCIL [18] Res18 $78.70$ $75.12$ $70.10$ $66.26$ $66.51$ $64.01$ $62.69$ $61.00$ $60.36$ $59.45$ $58.48$ $65.70$ $20.22$
GKEAL [66] Res18 $78.88$ $75.62$ $72.32$ $68.62$ $67.23$ $64.26$ $62.98$ $61.89$ $60.20$ $59.21$ $58.67$ $66.35$ $20.21$
NC-FSCIL [54] Res18 $80.45$ $75.98$ $72.30$ $70.28$ $68.17$ $65.16$ $64.43$ $63.25$ $60.66$ $60.01$ $59.44$ $67.28$ $21.01$
BiDistFSCIL [60] Res18 $79.12$ $75.37$ $72.80$ $69.05$ $67.53$ $65.12$ $64.00$ $63.51$ $61.87$ $61.47$ $60.93$ $67.34$ $18.19$
SAVC [39] Res18 $81.85$ $77.92$ $74.95$ $70.21$ $69.96$ $67.02$ $66.16$ $65.30$ $63.84$ $63.15$ $62.50$ $69.35$ $19.35$
F2M [36] Res18 $81.07$ $78.16$ $75.57$ $72.89$ $70.86$ $68.17$ $67.01$ $65.26$ $63.36$ $61.76$ $60.26$ $69.49$ $20.81$
CLIP (Bs.[32] ViT-B/16 $65.54$ $62.91$ $61.54$ $57.75$ $57.88$ $57.89$ $56.62$ $55.40$ $54.20$ $54.23$ $55.06$ $58.09$ 10.48
LP-DiF (Ours) ViT-B/16 83.94 80.59 79.17 74.30 73.89 73.44 71.60 70.81 69.08 68.74 68.53 74.00 $15.41$
Joint-LP (UB. of ours) ViT-B/16 83.9483.94{{83.94}}83.94 80.8380.83{{80.83}}80.83 79.4379.43{{79.43}}79.43 77.0677.06{{77.06}}77.06 76.3576.35{{76.35}}76.35 74.8974.89{{74.89}}74.89 73.6673.66{{73.66}}73.66 72.7972.79{{72.79}}72.79 71.8471.84{{71.84}}71.84 72.0672.06{{72.06}}72.06 71.8871.88{{71.88}}71.88 75.8875.88{{75.88}}75.88 12.0612.06{{12.06}}12.06
TABLE IV: Comparison between our LP-DiF and other replay-based FSCIL SOTA methods on mini-ImageNet. “Exemplar / cls” represents the number of exemplar of each class. Note that these existing FSCIL SOTA methods use the units of image as exemplar. “Disk Space / cls” represents the disk space consumed by the exemplar of each class.
Methods Exemplar / cls Disk Space / cls Avg
Replay [30] 1 image 51.19 KB 58.02
F2M [36] 5 images 255.95 KB 57.89
BDF [60] 1 image 51.19 KB 61.42
LP-DiF 2 vectors 1.15 KB 93.76
TABLE V: Comparison with state-of-the-art FSCIL methods on CIFAR-100. “Backbone.” represents the backbone of visual model.“Avg” represents the average accuracy of all sessions; the higher the value, the better performance. “PD” represents the Performance Drop rate; the lower the value, the better performance. “bs” is the abbreviation of “baseline”. “UB.” is the abbreviation of “Upper bound”.
Methods Backbone Accuracy in each session (%) $\uparrow$ Avg $\uparrow$ PD $\downarrow$
0 1 2 3 4 5  6 7 8
TOPIC [40] Res18 $64.10$ $55.88$ $47.07$ $45.16$ $40.11$ $36.38$ $33.96$ $31.55$ $29.37$ $42.62$ $34.73$
F2M [36] Res18 $64.71$ $62.05$ $59.01$ $55.58$ $52.55$ $49.96$ $48.08$ $46.28$ $44.67$ $53.65$ $20.04$
CEC [56] Res18 $73.07$ $68.88$ $65.26$ $61.19$ $58.09$ $55.57$ $53.22$ $51.34$ $49.14$ $59.53$ $23.93$
Replay [30] Res18 $74.40$ $70.20$ $66.54$ $62.51$ $59.71$ $56.58$ $54.52$ $52.39$ $50.14$ $60.78$ $24.26$
MetaFSCIL [12] Res18 $74.50$ $70.10$ $66.84$ $62.77$ $59.48$ $56.52$ $54.36$ $52.56$ $49.97$ $60.79$ $24.53$
GKEAL [66] Res18 $74.01$ $70.45$ $67.01$ $63.08$ $60.01$ $57.30$ $55.50$ $53.39$ $51.40$ $61.35$ $22.61$
C-FSCIL [21] Res12 $77.47$ $72.40$ $67.47$ $63.25$ $59.84$ $56.95$ $54.42$ $52.47$ $50.47$ $61.64$ $27.00$
FCIL [18] Res18 $77.12$ $72.42$ $68.31$ $64.47$ $61.18$ $58.17$ $56.06$ $54.19$ $52.02$ $62.66$ $25.10$
FACT [61] Res18 $78.22$ $72.40$ $68.57$ $64.73$ $61.40$ $58.57$ $56.30$ $53.83$ $51.72$ $62.86$ $26.50$
SAVC [39] Res18 $78.77$ $73.31$ $69.31$ $64.93$ $61.70$ $59.25$ $57.13$ $55.19$ $53.12$ $63.63$ $25.65$
BiDistFSCIL [60] Res18 $79.45$ $75.38$ $71.84$ $67.95$ $64.96$ $61.95$ $60.16$ $57.67$ $55.88$ $66.14$ $23.57$
NC-FSCIL [54] Res18 82.52 $76.82$ $73.34$ $69.68$ $66.19$ $62.85$ $60.96$ $59.02$ $56.11$ $67.50$ $26.41$
CLIP (Bs.[32] ViT-B/16 $74.44$ $72.96$ $72.21$ $70.49$ $70.18$ $70.00$ $69.81$ $69.23$ $68.37$ $70.86$ 6.07
LP-DiF (Ours) ViT-B/16 80.2380.23{80.23}80.23 77.75 76.78 74.62 74.03 73.87 73.84 72.96 72.02 75.12 $8.21$
Joint-LP (UB. of ours) ViT-B/16 80.2380.23{{80.23}}80.23 79.8579.85{{79.85}}79.85 78.6378.63{{78.63}}78.63 76.1376.13{{76.13}}76.13 75.3175.31{{75.31}}75.31 74.6774.67{{74.67}}74.67 74.2474.24{{74.24}}74.24 73.5873.58{{73.58}}73.58 73.3573.35{{73.35}}73.35 76.2276.22{{76.22}}76.22 6.886.88{{6.88}}6.88
Figure 4: Accuracy curves of our LP-DiF and comparison with counterparts on (a) SUN-397 and (b) CUB200*. our LP-DiF method significantly surpasses both CLIP and BiDistFSCIL, and attains performance levels that are very close to the respective upper bounds.
TABLE VI: Comparison with standard CIL methods based on pre-trained models on the three common benchmarks in terms of Avg. indicates our reproduction on FSCIL protocol.